OpenStudy (anonymous):

I need help with a 6th grade probability question. If three fair coins are tossed at the same time, what is the probability that 1 will show heads and two will show tails?

OpenStudy (anonymous):

Can anyone help with this?

OpenStudy (anonymous):

yes we can do this

OpenStudy (anonymous):

there are 8 possibilities when three coins are tossed, and since that is a fairly small number we can actually list them h, h, h h, h, t h, t , h t, h, h h, t, t t, h, t t, t, h t, t, t of these, 3 have one have one head and two tails, so your probability is \[\frac{3}{8}\]

OpenStudy (anonymous):

there is a snappier way to do this, but for 6th grade this should make the most sense

OpenStudy (anonymous):

I see that there are 4 possible outcomes if you are tossing them at the same time. That would eliminate the doubles of h,t,h and t,h,h and t,h,h and h,t,h. Right?

OpenStudy (anonymous):

"outcomes" in this case means all the possibilities for the heads or tails on the three coins. there are 2 for each coin (heads or tails) so by the "counting principle' there are \[2\times 2\times 2=8\] possibilities total. they are the ones i listed. hope that helps

OpenStudy (anonymous):

Thank you for the help. 3/8 is one of the answers.